OpenStudy (anonymous):

If f(x) = 5x2 − 4 and g(x) = x − 2, what is (f o g)(-3)? (A) -129 (B)1 (C)39 (D)121

OpenStudy (xapproachesinfinity):

\[f(g(x))=5(g(x))^2-4=2(x-2)^2-4=2[x^2-4x+4]-4=2x^2-8x+4\] so fog(-3)=2(-3)^2-8(-3)+4

OpenStudy (anonymous):

soo it would be 121

OpenStudy (xapproachesinfinity):

another way and perhaps easy is to plug -4 into g first g(-3)=-3-2=-5 then plug -5 into f f(-5)=?

OpenStudy (xapproachesinfinity):

that's not what i got

OpenStudy (anonymous):

what did u get

OpenStudy (xapproachesinfinity):

46

OpenStudy (anonymous):

i got it right it was 121

OpenStudy (xapproachesinfinity):

oh i did some mistake in calculations

OpenStudy (xapproachesinfinity):

i got 121 too just error lol

OpenStudy (xapproachesinfinity):

i used first one it should read 5x^2-20x+16

OpenStudy (xapproachesinfinity):

i by mistake multiplied by 2 there